OpenStudy (mikezack123):

The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 132 = x + 2(x - 9) What is the regular price of a child's ticket?

OpenStudy (whpalmer4):

The formula does not appear to match the words. The words say \[132 = 1* (\frac{1}{2})x + 4 *\frac{1}{2}(x-9) = \frac{1}{2}x + 2(x-9)\]

OpenStudy (mikezack123):

what do you mean?

OpenStudy (whpalmer4):

Well, your equation in the problem doesn't have the \(\frac{1}{2}\) in front of \(x\).

OpenStudy (mikezack123):

oh....

OpenStudy (mikezack123):

Thanks though! :)